OpenStudy (anonymous):

What is the answer to this quadratic equation of x^2-12x+22=0

OpenStudy (anonymous):

Do you know how to use the quadratic formula?

OpenStudy (anonymous):

could you please give me the answer to the equation, I would really appreciate it

OpenStudy (anonymous):

3+/- the square root of 2

OpenStudy (anonymous):

\[\left\{\left\{x\to 6-\sqrt{14}\right\},\left\{x\to 6+\sqrt{14}\right\}\right\}\to \{x\to 2.25834,x\to 9.74166\} \]

OpenStudy (anonymous):

A plot is attached.